OpenStudy (anonymous):

While rolling a wheel with a diameter of 19/2 ft makes 133 revolutions. Find the distance that is traveled by the wheel. Use 22/7 for pi.

OpenStudy (anonymous):

I'm unsure but might you just find the circumference and multiply by 133? I am unsure because I do not know the relationship between a wheels size and the distance it travels in a revolution

OpenStudy (anonymous):

Perimeter of the wheel: Pi*D = (22/7)*(19/2) = 418/14 = 29.86 ft Distance traveled by the wheel: 133 * 29.86 = ....

OpenStudy (anonymous):

actually the circumference is pi*r^2 not pi*d or pi*2r

OpenStudy (anonymous):

The area of a circle is : Pi*R^2. The perimeter is : Pi *D